OpenStudy (anonymous):

...I officially hate functions...the vertex of the function's graph is given. find c. y=c-x^2-2x (-1,3)

OpenStudy (anonymous):

You know that the point (-1,3) is a solution of the function. Plug in -1, and see what value of c will work to make it equal to 3.

OpenStudy (anonymous):

do you really have nothing better to do?

OpenStudy (anonymous):

im gonna be honest, i really dont LOL

OpenStudy (anonymous):

well then i feel sorry for you

OpenStudy (anonymous):

Ignoring that, does what I suggested make any sense? We know that WHEN x is -1, y = 3. So we plug in -1 and set equal to 3, like this c-(-1)^2-2(-1) = 3

OpenStudy (anonymous):

ya it makes sense...i'm just working through my options at this point...

OpenStudy (anonymous):

Haha okay, well, once you've done that, the rest is just simple algebra.

OpenStudy (anonymous):

yes it is:) i believe my final answer is c=2....thanks once again lol :)

OpenStudy (anonymous):

c-(-1)^2-2(-1) = 3 c-1-(-2) = 3 c-1+2 = 3 c+1 = 3 c = 2 Very good =)